Show all work and solve
y = x2 + 3
y = x + 5

Answer by ankor@dixie-net.com(22740)

y = x^2 + 3
y = x + 5
y = y; therefore
x^2 + 3 = x + 5
x^2 - x + 3 - 5 = 0
x^2 - x - 2 = 0
Factors to
(x - 2)(x + 1) = 0
Two solutions
x = 2
x = -1
:
You can check both solutions in both equations